Consider the differential equation (1-x²)y - 5xy - 3 y = 0 1. Find its general solution y = Xar, x in the form y = doy1(x)
Answer #1

Date Daga Do given differential equation (i- de 2) y - sxy - 3y = 0 let y = Eanxh y(x) = Enana y(x) = n(n-1) an xn-2 ho hi000 Page -34 5 han -n(ny) an +n+2) Chti) au +2 m-n-sn-3) + (n+2) (h+I) 2 -30 > anta (n+3) Cht) Fo+2) (+1) h+2 so IN प m:2 5_9lim x2 2 (n+1) Radius of Convergence 4 12) = E 2 int! Gn+1) cant) 3.1 2nt3 2n + 3 L 121 Radius of Cony lim ht an+2) (anti) (